OpenStudy (anonymous):

Square root of ((-6)^5)^4) =

OpenStudy (anonymous):

\[\sqrt{((-6)^5)^4}\]

OpenStudy (anonymous):

((-6)^5)^4=6^20 so sqrt(6^20) = (6^20)^(1/2) = 6^10

OpenStudy (anonymous):

Thank youuuuu =)